OpenStudy (anonymous):

Quick question just for clarification: |-11| = -11?

ganeshie8 (ganeshie8):

Nope. Absolute value = distance distance can never be negative

OpenStudy (anonymous):

Ok so in the situation |2x+3| = x -5, the answer would be no solution correct?

ganeshie8 (ganeshie8):

why

OpenStudy (anonymous):

I ended up with the answers |-13| = -13 and |13/3| = -13/3

OpenStudy (anonymous):

and the two answer choices i'm debating is {-8, 2/3} or No solution

ganeshie8 (ganeshie8):

then maybe you're right. no solutions.

OpenStudy (anonymous):

Gracias!

ganeshie8 (ganeshie8):

gracias :)